# Université Paris-Saclay, ENS Paris-Saclay, CNRS, LMF

### v1

@@ 1 Introduction
Third, we have defined uniform interpolants with respect to an entailment relation $\vdash$, but there exists a subtly different definition in the literature. Indeed, let $\to$ be a binary symbol in the language under consideration. One may then also reasonably define [^2] For instance, this is how uniform interpolation is defined in Definition 24 in Section 3.3 of [32]. implication-based uniform interpolants by replacing, in Definition 1.2, all assertions of the form $\alpha\vdash\beta$’ by $\top\vdash\alpha\to\beta$’, where $\top$ is a fixed constant symbol in the language, suitably connected to the operation $\to$. Let us say that the entailment relation $\vdash$ has the deduction property if, for all formulas $\alpha$ and $\beta$, $\alpha\vdash\beta$ is equivalent to $\top\vdash\alpha\to\beta$. If the entailment relation $\vdash$ does not have this deduction property, then the notions of uniform interpolant in the sense of Definition 1.2 and the implication-based uniform interpolant are not the same in general. Indeed, when $\vdash$ is the global consequence relation for the modal logic $\mathbf{K}$, uniform interpolants in the sense of Definition 1.2 do not always exist, while implication-based uniform interpolants do exist [23, Sec. 6]; we come back to this in Remark 2.16 below. In light of this, the uniform interpolants of Definition 1.2 are sometimes called deductive uniform interpolants for emphasis.
@@
Proof methods for uniform interpolation can be divided, roughly, into two strands: one is syntactic and relies on the existence of a well-behaved sequent calculus for the entailment relation $\vdash$. This is how uniform interpolation was established for intuitionistic entailment in Pitts’ original proof [60]. Pitts’ proof relies on a strongly terminating proof calculus for intuitionistic propositional logic, and has been the subject of intensive proof-theoretic investigations and extensions to other logics; see, for instance, Section 3.3 in [32] and the references therein; we will also make some more remarks on this method in Section 5 of this chapter. The other proof method, which will be our focus in this chapter, starts from a semantic characterization of the entailment relation $\vdash$, typically in terms of a consequence relation on Kripke models. One then shows that bisimulation quantifiers are definable in the propositional language, and deduces from this the existence of implication-based uniform interpolants. This approach to proving implication-based uniform interpolation was developed soon after [60] as an alternative proof for Theorem 1.1 [27], and was also shown to apply to modal logics such as Gödel-Löb logic $\mathbf{GL}$ [68, Cor. 2.13], modal logic $\mathbf{K}$ [23, Thm 6.1] and Grzegorczyk logic $\mathbf{S4Grz}$ [72, Sec. 8].
@@
Note that our focus in this chapter is on uniform interpolation for propositional logics. In the classical setting, first order logic does not have uniform interpolants, see for instance Theorem 2.18 in [13]. There are substructural predicate logics with uniform interpolation [4]. We do not know at present whether or not intuitionistic predicate logic has uniform interpolation.
@@ 2.1 Models and Bisimilarity
Intuitionistic entailment is sound and complete with respect to finite Kripke models. More precisely, if $\phi(\overline{p})$ and $\psi(\overline{p})$ are formulas, then the intuitionistic entailment $\phi\vdash\psi$ holds if, and only if, for any finite pointed $\overline{p}$-model $(M,w)$, if $M,w\Vdash\phi$ then $M,w\Vdash\psi$. For a proof of this foundational equivalence, see, e.g., [11, Thm. 2.43]. In light of this equivalence, when $\phi(\overline{p})$ is a formula and $\overline{q}$ is any set of variables with $\overline{p}\subseteq\overline{q}$, we define
@@ 2.4 Extensions of the Semantic Method to Other Logics
The difference between implication-based and deductive uniform interpolants can be seen clearly in the case of the modal logic $\mathbf{K}$. An algorithm for computing implication-based uniform interpolants for the logic $\mathbf{K}$ is given in [8], and is also outlined in Section 3.3 of [32]. The fact that deductive uniform interpolants do not always exist in $\mathbf{K}$ is essentially due to [26, Thm. 37], where a counterexample is given in the syntax of description logic. The failure is closely related to the failure of the deduction theorem for the global consequence relation, for which see, e.g., [6, Sec. 1.5]. We here give a slightly simpler example, due to G. Metcalfe (personal communication); essentially the same example can be used to show that any strongly complete modal logic which has right uniform interpolants must be weakly transitive. Here, one may even weaken the assumption of having right uniform interpolants to being coherent in the sense of Theorem 3.19 below [45, Thm. 4.1].
@@ 3.1 Uniform Interpolants as Adjoints
Let $\mathcal{V}$ be a variety of algebras, in the sense of universal algebra, see, e.g., Section 2 of [56]. We write $\mathbf{F}_{\mathcal{V}}(\overline{x})$ for the free $\mathcal{V}$-algebra over a set of variables $\overline{x}$. Recall that, in Section 3 of that chapter, a consequence relation $\vdash_{\mathcal{V}}$ was associated to $\mathcal{V}$, and it was shown in Lemma 9 of the same chapter that, for any set of equations $\Sigma\cup\{E\}$ with variables in a set $\overline{x}$, we have $\Sigma\vdash_{\mathcal{V}}E$ if, and only if, the equation $E$ belongs to the congruence on $\mathbf{F}_{\mathcal{V}}(\overline{x})$ generated by $\Sigma$. For any homomorphism $h\colon A\to B$ between algebras in $\mathcal{V}$, we have an adjoint pair of maps
@@ 4 Model-complete Theories from Uniform Interpolants
The starting point of the work of S. Ghilardi and M. Zawadowski in [29, 30] is that Pitts’ Theorem 1.1 can be used to establish that the first order theory of Heyting algebras has a model completion [29, p. 29]. The concept of model completion that is used here originates in A. Robinson’s model-theoretic algebra [65, 66], and closely relates to the idea of quantifier elimination. The prime example is the theory of algebraically closed fields, which is the model completion of the first order theory of integral domains. An example closer to the kind we want to consider here is the first order theory of Boolean algebras, which has a model completion, namely, the theory of atomless Boolean algebras, see, e.g., [63, Thm. 6.21]. We now recall the definitions and some basic facts about these concepts, referring to, e.g. [12, Sec. 3.5] for more details.
@@
It is well-known that the class of Heyting algebras has the amalgamation property, see for instance [54, Prop. 1] or [20] for early algebraic proofs of this fact. In light of the above remarks, it therefore suffices to show that there is a first order theory whose models are exactly the existentially closed Heyting algebras.
@@
where we use the uniform interpolants for propositional formulas, which exist according to Theorem 1.1. Using the definition of uniform interpolants, one may then prove that, for any Heyting algebra $A$ and $\overline{a}\in A^{\overline{x}}$, we have $A\models\Psi^{\prime}(\overline{x}/\overline{a})$ if, and only if, $A[y]/\alpha\models\Psi(\overline{x}/\overline{a})$, where $A[y]$ denotes the coproduct of $A$ with the free Heyting algebra over $y$, and $\alpha$ is the congruence generated by the pairs $(\phi_{i}(\overline{x}/\overline{a},y),\top)$, for $i=1,\dots,n$. Therefore, a Heyting algebra $A$ satisfies the existential closedness property for the formula $\Psi(\overline{x},y)$ and a tuple $\overline{a}\in A^{\overline{x}}$ if, and only if, $A\models(\Psi^{\prime}\rightarrow\Psi)(\overline{x}/\overline{a})$. Thus, an existentially closed Heyting algebra must satisfy the first order sentence $\forall\overline{x}.(\Psi^{\prime}\rightarrow\Psi)$. This is the essential property required for axiomatizing existentially closed Heyting algebras. Using this property, a general argument (see, e.g., [12, Thm. 3.5.20]) shows that the model completion of $\mathrm{Th}_{\mathsf{FO}}(\mathcal{H})$ is obtained by adding to the theory $\mathrm{Th}_{\mathsf{FO}}(\mathcal{H})$, for any choice of $\overline{x},y,\phi_{1},\dots,\phi_{n},\psi_{1},\dots,\psi_{m}$ as above, the formula $\forall\overline{x}.(\Psi^{\prime}\to\Psi)$.
@@ 5 Syntactic Methods, Further Results and Connections
Our focus in this chapter so far has been on connections between uniform interpolation and semantics. The syntactic method of establishing uniform interpolation is discussed in more detail in Section 3 of [32] in this volume; we will here just make some brief remarks about the method and its relation to the semantic method.
@@
The syntactic method was in particular the one that was originally used by Pitts to establish Theorem 1.1 in [60], and was subsequently used for proving the implication-based uniform interpolation property for a great number of other logics; for instance, this was done by Bílková [8] for the basic modal logic $\mathbf{K}$ and for the modal logic of reflexive frames, $\mathbf{T}$, and for provability logics $\mathbf{GL}$ and $\mathbf{Grz}$ in [7], also see [9]. The method was further extended by Iemhoff [36, 37], who identifies precise general criteria for sequent calculi that allow one to establish uniform interpolation; these ideas were subsequently applied to obtain syntactic proofs of uniform interpolation for a number of intuitionistic modal logics, see, e.g., [31, 3, 69], and the further references given in [32] of this volume. In the context of substructural logic, Alizadeh, Derakhshan, and Ono [4] establish uniform interpolation for full Lambek calculus and for linear and affine logics without exponentials. Interestingly, in that context, even the predicate versions of the logics have uniform interpolation, in contrast with the case of classical logic, see e.g. Theorem 2.17 in [13].
@@
In Table 1, we summarize the current state of the art of the literature on uniform interpolation for a number of common logics, and the existence of model companions for their corresponding classes of algebras. Note that this table also implies a number of open questions: for all the empty cells, at the time of writing, the author is unaware of any answers, yet hopeful that readers of this chapter will be inspired to produce some. [^8] Just before this chapter was going to press in December 2025, the author was informed that uniform interpolation for iGL has been proved by I. van der Giessen, B. Sierra Miranda, and G. Menéndez Turata. For more information, we refer to the table maintained by T. Kurahashi at https://www2.kobe-u.ac.jp/˜tk/jp/notes/ULIP.html, which surveys Craig interpolation, implication-style uniform interpolation, and their Lyndon variants, for a large number of intermediate, modal, and intuitionistic modal logics.

### v2

@@ 1 Introduction
Third, we have defined uniform interpolants with respect to an entailment relation $\vdash$, but there exists a subtly different definition in the literature. Indeed, let $\to$ be a binary symbol in the language under consideration. One may then also reasonably define [^2] For instance, this is how uniform interpolation is defined in Definition 24 in Section 3.3 of [32]. hard-coded references in footnote implication-based uniform interpolants by replacing, in Definition 1.2, all assertions of the form $\alpha\vdash\beta$’ by $\top\vdash\alpha\to\beta$’, where $\top$ is a fixed constant symbol in the language, suitably connected to the operation $\to$. Let us say that the entailment relation $\vdash$ has the deduction property if, for all formulas $\alpha$ and $\beta$, $\alpha\vdash\beta$ is equivalent to $\top\vdash\alpha\to\beta$. If the entailment relation $\vdash$ does not have this deduction property, then the notions of uniform interpolant in the sense of Definition 1.2 and the implication-based uniform interpolant are not the same in general. Indeed, when $\vdash$ is the global consequence relation for the modal logic $\mathbf{K}$, uniform interpolants in the sense of Definition 1.2 do not always exist, while implication-based uniform interpolants do exist [23, Sec. 6]; we come back to this in Remark 2.16 below. In light of this, the uniform interpolants of Definition 1.2 are sometimes called deductive uniform interpolants for emphasis.
@@
Proof methods for uniform interpolation can be divided, roughly, into two strands: one is syntactic and relies on the existence of a well-behaved sequent calculus for the entailment relation $\vdash$. This is how uniform interpolation was established for intuitionistic entailment in Pitts’ original proof [60]. Pitts’ proof relies on a strongly terminating proof calculus for intuitionistic propositional logic, and has been the subject of intensive proof-theoretic investigations and extensions to other logics; see, for instance, Section 3.3hard-coded in [32] and the references therein; we will also make some more remarks on this method in Section 5 of this chapter. The other proof method, which will be our focus in this chapter, starts from a semantic characterization of the entailment relation $\vdash$, typically in terms of a consequence relation on Kripke models. One then shows that bisimulation quantifiers are definable in the propositional language, and deduces from this the existence of implication-based uniform interpolants. This approach to proving implication-based uniform interpolation was developed soon after [60] as an alternative proof for Theorem 1.1 [27], and was also shown to apply to modal logics such as Gödel-Löb logic $\mathbf{GL}$ [68, Cor. 2.13], modal logic $\mathbf{K}$ [23, Thm 6.1] and Grzegorczyk logic $\mathbf{S4Grz}$ [72, Sec. 8].
@@
Note that our focus in this chapter is on uniform interpolation for propositional logics. In the classical setting, first order logic does not have uniform interpolants, see for instance Theorem 2.17hard-coded in [10]. There are substructural predicate logics with uniform interpolation [4]. We do not know at present whether or not intuitionistic predicate logic has uniform interpolation.
@@ 2.1 Models and Bisimilarity
Intuitionistic entailment is sound and complete with respect to finite Kripke models. More precisely, if $\phi(\overline{p})$ and $\psi(\overline{p})$ are formulas, then the intuitionistic entailment $\phi\vdash\psi$ holds if, and only if, for any finite pointed $\overline{p}$-model $(M,w)$, if $M,w\Vdash\phi$ then $M,w\Vdash\psi$. For a proof of this foundational equivalence, see, e.g., [12, Thm. 2.43]. In light of this equivalence, when $\phi(\overline{p})$ is a formula and $\overline{q}$ is any set of variables with $\overline{p}\subseteq\overline{q}$, we define
@@ 2.4 Extensions of the Semantic Method to Other Logics
The difference between implication-based and deductive uniform interpolants can be seen clearly in the case of the modal logic $\mathbf{K}$. An algorithm for computing implication-based uniform interpolants for the logic $\mathbf{K}$ is given in [8], and is also outlined in Subsection 3.3hard-coded of [32]. The fact that deductive uniform interpolants do not always exist in $\mathbf{K}$ is essentially due to [26, Thm. 37], where a counterexample is given in the syntax of description logic. The failure is closely related to the failure of the deduction theorem for the global consequence relation, for which see, e.g., [6, Sec. 1.5]. We here give a slightly simpler example, due to G. Metcalfe (personal communication); essentially the same example can be used to show that any strongly complete modal logic which has right uniform interpolants must be weakly transitive. Here, one may even weaken the assumption of having right uniform interpolants to being coherent in the sense of Theorem 3.19 below [45, Thm. 4.1].
@@ 3.1 Uniform Interpolants as Adjoints
Let $\mathcal{V}$ be a variety of algebras, in the sense of universal algebra, see, e.g., Section 2hard-coded of [56]. We write $\mathbf{F}_{\mathcal{V}}(\overline{x})$ for the free $\mathcal{V}$-algebra over a set of variables $\overline{x}$. Recall that, in Section 3hard-coded of that chapter, a consequence relation $\vdash_{\mathcal{V}}$ was associated to $\mathcal{V}$, and it was shown in Lemma 9hard-coded of the same chapter that, for any set of equations $\Sigma\cup\{E\}$ with variables in a set $\overline{x}$, we have $\Sigma\vdash_{\mathcal{V}}E$ if, and only if, the equation $E$ belongs to the congruence on $\mathbf{F}_{\mathcal{V}}(\overline{x})$ generated by $\Sigma$. For any homomorphism $h\colon A\to B$ between algebras in $\mathcal{V}$, we have an adjoint pair of maps
@@ 4 Model-complete Theories from Uniform Interpolants
The starting point of the work of S. Ghilardi and M. Zawadowski in [29, 30] is that Pitts’ Theorem 1.1 can be used to establish that the first order theory of Heyting algebras has a model completion [29, p. 29]. The concept of model completion that is used here originates in A. Robinson’s model-theoretic algebra [65, 66], and closely relates to the idea of quantifier elimination. The prime example is the theory of algebraically closed fields, which is the model completion of the first order theory of integral domains. An example closer to the kind we want to consider here is the first order theory of Boolean algebras, which has a model completion, namely, the theory of atomless Boolean algebras, see, e.g., [63, Thm. 6.21]. We now recall the definitions and some basic facts about these concepts, referring to, e.g. [13, Sec. 3.5] for more details.
@@
It is well-known that the class of Heyting algebras has the amalgamation property, see for instance [53, Prop. 1] or [20] for early algebraic proofs of this fact. In light of the above remarks, it therefore suffices to show that there is a first order theory whose models are exactly the existentially closed Heyting algebras.
@@
where we use the uniform interpolants for propositional formulas, which exist according to Theorem 1.1. Using the definition of uniform interpolants, one may then prove that, for any Heyting algebra $A$ and $\overline{a}\in A^{\overline{x}}$, we have $A\models\Psi^{\prime}(\overline{x}/\overline{a})$ if, and only if, $A[y]/\alpha\models\Psi(\overline{x}/\overline{a})$, where $A[y]$ denotes the coproduct of $A$ with the free Heyting algebra over $y$, and $\alpha$ is the congruence generated by the pairs $(\phi_{i}(\overline{x}/\overline{a},y),\top)$, for $i=1,\dots,n$. Therefore, a Heyting algebra $A$ satisfies the existential closedness property for the formula $\Psi(\overline{x},y)$ and a tuple $\overline{a}\in A^{\overline{x}}$ if, and only if, $A\models(\Psi^{\prime}\rightarrow\Psi)(\overline{x}/\overline{a})$. Thus, an existentially closed Heyting algebra must satisfy the first order sentence $\forall\overline{x}.(\Psi^{\prime}\rightarrow\Psi)$. This is the essential property required for axiomatizing existentially closed Heyting algebras. Using this property, a general argument (see, e.g., [13, Thm. 3.5.20]) shows that the model completion of $\mathrm{Th}_{\mathsf{FO}}(\mathcal{H})$ is obtained by adding to the theory $\mathrm{Th}_{\mathsf{FO}}(\mathcal{H})$, for any choice of $\overline{x},y,\phi_{1},\dots,\phi_{n},\psi_{1},\dots,\psi_{m}$ as above, the formula $\forall\overline{x}.(\Psi^{\prime}\to\Psi)$.
@@ 5 Syntactic Methods, Further Results and Connections
Our focus in this chapter so far has been on connections between uniform interpolation and semantics. The syntactic method of establishing uniform interpolation is discussed in more detail in Section 3 hard-coded of [32] in this volume; we will here just make some brief remarks about the method and its relation to the semantic method.
@@
The syntactic method was in particular the one that was originally used by Pitts to establish Theorem 1.1 in [60], and was subsequently used for proving the implication-based uniform interpolation property for a great number of other logics; for instance, this was done by Bílková [8] for the basic modal logic $\mathbf{K}$ and for the modal logic of reflexive frames, $\mathbf{T}$, and for provability logics $\mathbf{GL}$ and $\mathbf{Grz}$ in [7], also see [9]. The method was further extended by Iemhoff [36, 37], who identifies precise general criteria for sequent calculi that allow one to establish uniform interpolation; these ideas were subsequently applied to obtain syntactic proofs of uniform interpolation for a number of intuitionistic modal logics, see, e.g., [31, 3, 69], and the further references given in [32] of this volume. In the context of substructural logic, Alizadeh, Derakhshan, and Ono [4] establish uniform interpolation for full Lambek calculus and for linear and affine logics without exponentials. Interestingly, in that context, even the predicate versions of the logics have uniform interpolation, in contrast with the case of classical logic, see e.g. Theorem 2.17 hardcoded in [10].
@@
In Table 1, we summarize the current state of the art of the literature on uniform interpolation for a number of common logics, and the existence of model companions for their corresponding classes of algebras. Note that this table also implies a number of open questions: for all the empty cells, at the time of writing, the author is unaware of any answers, yet hopeful that readers of this chapter will be inspired to produce some. [^8] Just before this chapter was going to press in December 2025, the author was informed that uniform interpolation for iGL has been proved by I. van der Giessen, B. Sierra Miranda, and G. Menéndez Turata. For more information, we refer to the table maintained by T. Kurahashi on the webpage https://www2.kobe-u.ac.jp/˜tk/jp/notes/ULIP.html, which surveys Craig interpolation, implication-style uniform interpolation, and their Lyndon variants, for a large number of intermediate, modal, and intuitionistic modal logics.